OpenStudy (anonymous):

medal Find the sum of the summation of 2 i minus 9, from i equals 3 to 10

OpenStudy (anonymous):

answer choices 52 26 13 32

OpenStudy (anonymous):

so u need help?

OpenStudy (anonymous):

there

OpenStudy (anonymous):

what

OpenStudy (ybarrap):

Is this the problem? $$ \Large{ \sum_{i=3}^{10}\left ( 2i-9\right ) } $$ If so, let's rearrange indices so that i starts at zero and simplify: $$ \Large{ \sum_{i=0}^{7}\left ( 2(i+3)-9\right )=\sum_{i=0}^{7}\left ( 2i+6-9\right )\\ =\sum_{i=0}^{7}\left ( 2i-3\right )=\left (2\sum_{i=0}^{7}i\right )-3\times 8 } $$ Does this make sense? Can you finish?

OpenStudy (anonymous):

32